Getler Inc.'s projected capital budget is $2,000,000, its target capital structure is 40% debt and 60% equity, and its forecasted net income is $1,000,000. If the company follows a residual dividend policy, how much dividends will it pay or, alternatively, how much new stock must it issue?       Dividends     Stock Issued

 

Select one:

a. $514,425     $162,901 b. $541,500     $171,475 c. $570,000     $180,500 d. $600,000     $190,000 e. $          0     $200,000
Answer #1

okay so our company follows a Residual Dividend policy. First Let's understand what is this policy?

Residual dividend policy is something used by companies which finance their capital requirements through money that is generated in house.That is they would pay dividends only when all the capital requirements have been met through the use of internal money rather than going to market.

Our situation;

Capital Budget - $ 2,000,000 Capital structure is 40% Debt and 60% Equity

So lets calculate our Forecasted Debt and Equity amount first:

Debt @ 40% = 2,000,000 * 40% = $800,000

Equity @ 60% = 2,000,000 * 60% = $1,2000,000

Our income is $ 1,000,000

Now we can use our internal income of 1,000,000 to finance our equity requirement of 1,2000,000

Residual amount = 1,000,000 - 1,2000,000 = -200,000

That means our company needs to issue $ 200,000 worth of stock more and it won't be able to pay any dividends as yet.

Our answer option E) 0 $ 200,000